OpenStudy (anonymous):

find the circumference of a circle with a radius of 4cm use 3.14 for pi

OpenStudy (saifoo.khan):

Do u know the formula?

OpenStudy (saifoo.khan):

Please Try first.

OpenStudy (anonymous):

8x3.14=25.12

OpenStudy (anonymous):

ummm pi r 2???

OpenStudy (saifoo.khan):

C = pi D. D = diameter.

OpenStudy (anonymous):

circumference = 2*pi*r = 2*3.14*4cm = 8*3.14cm = 25.12 cm see here u just want to know the formula then after that it is much easier

OpenStudy (anonymous):

or pi r 2

OpenStudy (saifoo.khan):

Right, same thing.

OpenStudy (anonymous):

yeah okay so 3.14 x 8

OpenStudy (saifoo.khan):

Correct.

OpenStudy (saifoo.khan):

Now i will give u a Good Answer when u will try yourself too! :)

OpenStudy (anonymous):

yay is that how its aligned though too?

OpenStudy (saifoo.khan):

C = 2 * pi * r C = 3.142 * 8 C = 25.136cm.

OpenStudy (anonymous):

thanks :D

OpenStudy (saifoo.khan):

No Problem!

OpenStudy (anonymous):

Dude, I said the answer 10 hours ago.

OpenStudy (anonymous):

wait i got 25.12 saifoo.khan

OpenStudy (saifoo.khan):

That is same thing. Decimals dont matter much if the question doesn't says about worrying for decimals. u are getting 2 in the end coz u took 3.14 for pi. I took 3.142, where 3 decimal places are used as a standard.

OpenStudy (anonymous):

but it does matter in my math book

OpenStudy (anonymous):

shanny u got 25.12 because u took pi as 3.14 and saifoo took pi as 3.142

OpenStudy (anonymous):

shanny u got 25.12 because u took pi as 3.14 and saifoo took pi as 3.142

OpenStudy (saifoo.khan):

my answer is also correct. Your is also correct/

OpenStudy (anonymous):

shanny u got 25.12 because u took pi as 3.14 and saifoo took pi as 3.142

OpenStudy (anonymous):

ohhhhh kk thanks :D

OpenStudy (anonymous):

shanny u got 25.12 because u took pi as 3.14 and saifoo took pi as 3.142

OpenStudy (anonymous):

shanny u got 25.12 because u took pi as 3.14 and saifoo took pi as 3.142

OpenStudy (anonymous):

shanny u got 25.12 because u took pi as 3.14 and saifoo took pi as 3.142

OpenStudy (anonymous):

shanny u got 25.12 because u took pi as 3.14 and saifoo took pi as 3.142

OpenStudy (anonymous):

shanny u got 25.12 because u took pi as 3.14 and saifoo took pi as 3.142

OpenStudy (anonymous):

shanny u got 25.12 because u took pi as 3.14 and saifoo took pi as 3.142

OpenStudy (anonymous):

oh sorry for too much cmments

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

yeah